Freshly dissected cortical bone contains ~10 vol.%water. Dried cortical bone contains ~70 wt.% minerals and ~30 wt.% collagen proteins. The density of carbonated hydroxyapatite is ~2.8 g/cm3and the density of dried collagen is ~1.3 g/cm3 .

a.Calculate the volume percentage of minerals and proteins (each) in dehydrated bone.

b.Calculate the volume percentage of minerals and proteins (each) in hydrated bone.
Answer #1

A )mineral wt in dry bone is 70% w/w

70 gm in 100gm of dry bone

Density of hydroxyapatite 2.8g/cm3= 70gm/vol

Vol= 70/2.8=25cm3 or 25ml

Protein wt 30% means 30gm in 100gm of dry bone

Density = 1.3g/cm3= 30gm/vol

Vol= 30/1.3 =23.07 cm3= 23.07ml

Hope it's clear..thanks